OpenStudy (anonymous):

A balloonist happens to drop his metal ballpoint pen from his balloon as he is taking notes on his flight. Because his pen has a streamlined shape, air resistance has little effect on its fall. The pen strikes the ground at 22 m/s. How high was the balloon? (Points : 6) 25 m 49 m 850 m 9,500 m

OpenStudy (anonymous):

u have to find acceleration to find height

OpenStudy (anonymous):

the formula that can be used is v^2=u^2+2as

OpenStudy (anonymous):

wht is your answer

OpenStudy (anonymous):

\(\vec v_f^2=\vec v_i^2+2\vec a \cdot \vec S\) since \(\vec v_i=0\hat j\),\(\vec v_f=-22\hat j\),\(\vec a=-10\hat j \), plug it in to get, \((-22\hat j)^2=2(-10\hat j) \cdot \vec S\) \(\vec S=-24.2\hat j\) do you need further help?

OpenStudy (anonymous):

@Shadowys u could've aloud @megen505 to solve this q' her own and understand. at least a chance